OpenStudy (anonymous):

Solve for x. 6 – 2x – x = 18 A. –24 B. –12 C. –8 D. –4

OpenStudy (anonymous):

Please Help Will Give Medal

OpenStudy (anonymous):

First add the x's to get -x. Then subtract 6 from both sides. Let me know what you get.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Tanks for fanning me.

OpenStudy (anonymous):

Can you draw kinda how to do it

OpenStudy (anonymous):

\[6-2x-x=18\] \[6-x=18\] \[-6 -6\] \[-x=\]

OpenStudy (anonymous):

It's sutract 6 from both sides meaning you should subtract 6 from 18.

OpenStudy (anonymous):

so 12

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

No just 12. Good Job!!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

that,s what i though it was before i got on here Thanks

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

That's

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

I got what you were saying. Close this question for me.

OpenStudy (anonymous):

k